Please help, I have re worked this problem and I keep getting mixed solutions.Please write clearly Extra Credit: There will be no partial credit. Answers must be correct. (10 pts) Show how you would accomplish the following
Answer #1

The given tertiary alcohol is OH 4-Methylheptan-4-ol This is a tertiary alcohol To prepare this we can start with an ester 2Preparation opestry from Ethanol: oxidation o Cro3 он RoH + Ho! - H₂o Ethyl Acetate Preparation op propyl magnesiumbromide fr0 = c =CH-CH3. C - CH₂ - CH₂ Reduction LiAlHa Bs 83 CH₂ - CH₂ - CH₂-OH (o=P-BY) mg, Etzo 2+ Mg Br = Mg Bro finally i slave. 4